OpenStudy (anonymous):

Solve for x. lnx - ln(x +1) = 1

OpenStudy (anonymous):

lnx(1/x+1)=1

OpenStudy (anonymous):

LN properties subtraction is really division

OpenStudy (lalaly):

ln (x/x+1) = 1 x/x+1 = e^1 x=e^1 +xe^1 x-xe^1 = e^1 x(1-e^1)=e^1 x=[e^1/(1-e^1)]

OpenStudy (anonymous):

apparently though, e/1-e is extraneous so its no solution, i don't understand that...

OpenStudy (anonymous):

to remove a ln you must raise to the power of e

OpenStudy (anonymous):

you cannot take the ln of a number less then 0

OpenStudy (anonymous):

i don't see how that applies to this though..